Space Stations

NASA ISS On-Orbit Status 26 May 2014

By Marc Boucher
Status Report
May 27, 2014
Filed under , ,
NASA ISS On-Orbit Status 26 May 2014
NASA ISS On-Orbit Status 26 May 2014.

Today: European Physiology Module (EPM) Rack Reconfiguration: Commander (CDR) Swanson reconfigured the EPM rack in preparation for the Plasma-Kristall 4 investigation planned for Increment 41. He swapped out close-out panels, stowage containers, and CARDIOLAB within the EPM rack.
Remote Power Control Module (RPCM) Trip: Early this morning, RPCM 4B_A RPC #4 tripped. This Remote Power Controller (RPC) feeds the Loop A Early External Thermal Control System (EETCS) pump. While this pump is plumbed into PVTCS 2B as part of the hybrid configuration due to the PVTCS leak, the pump was not running. The pump has been powered to provide quantity sensor data. Ground controllers have reviewed the trip data and agree this is a real overcurrent event. There is no immediate impact to cooling as the PVTCS 2B Loop continues to operate nominally in the hybrid configuration. A Timeliner bundle is currently monitoring 6 sensors on the Loop B leg, two of which come from this pump. With the RPC open the Timeliner will continue to see the last value from the accumulator quantities for the Loop A pump. Since this won’t provide accurate insight to a future leak SPARTAN has inhibited the Timeliner bundle and re-enabled the 2B PFCS FDIR.

Camera Light Pan Tilt Assembly (CLPA) Remove & Replace (R&R) Preparations: Following last Friday night’s operations when ground controllers removed the Space Station Remote Manipulator System (SSRMS) Boom B faulty CLPA, today Swanson installed a new CLPA on the Japanese Experiment Module (JEM) Airlock slide table. Tomorrow the Special Purpose Dexterous Manipulator (SPDM) will be used to install the CLPA on the SSRMS End-B.

Lamp Housing Assembly (LHA) R&R: Swanson performed two LHA R&Rs, both in Node 1. Swanson successfully replaced the lamps in the Overhead-Starboard-2 and Port-Deck-2 locations (NOD1OS2 and NOD1PD2). Replacement of these two LHAs returns the Node 1 module to fully operational lighting conditions.

Today’s Planned Tasks

RSS1,2 Reboot [Complete]
SM ??? (Caution & Warning Panel) Test / Manual Controls [???], [Complete]
Laptop RS1(2) Reboot / Onboard Computer System (???), [Complete]
???-?? ??-? connector inspection behind pan. [Complete]
Calf volume measurement / ?? volume [Complete]
SEISMOPROGNOZ. Data copying from ???? HDD to external media (start). r/g 5701 [Complete]
On MCC Go ?1 ??? cartridge regeneration (start) [Complete]
DPC (S-band) [Complete]
Acoustic Dosimeter Preparation [Complete]
S/G1 Voice Loop Comm. Check with ??? Baikonur (S-band) [Complete]
SM ISS ???? “??? Z0000? removal. r/g 5695 [Complete]
Antivirus Database update on Auxiliary Computer Systems (???) laptops / r/g 5705 [Complete]
G1-camcorder installation [Complete]
JEM Airlock folding table extension [Complete]
CLPA installation in JEM [Complete]
??1, ??2 dust collector filters R&R and ?1, ?2 fan screens cleaning in DC1 [Complete]
JEM Airlock table folding [Complete]
Fan B3 screen cleaning in DC1 [Complete]
JEM Airlock – pressure release [Complete]
G1-camcorder close-out ops [Complete]
Humidity absorber retrieval from MERLIN. Kit replacement [Complete]
EPM – rack cleaning [Complete]
OF GAGARIN FROM SPACE. KENWOOD radio station settings restore. [Complete]
On MCC Go Re-mate cables to ???2 from PCE (????) to ?????. [Complete]
Restoring S? interior into initial configuration after dismantling PCE (????) equipment r/g 5695 [Complete]
MATRYOSHKA-R. Monitoring Liulin-5 equipment readings / r/g 4110 [Complete]
JEM Airlock – leak check [Complete]
TV Comm. (???) test with multicast mpeg2 flow transmission using ESA MPEG2 ENCODER prior to Soyuz #713 docking. [Complete]
SM Rodnik system tank valves (??, ?? and ??) scheduled maintenance [Complete]
VEG-daily visual inspection of plants [Complete]
MELFI2 – placing ice cubes into MELFI 2 [Complete]
Filling (separation) of ??? (???) for “Elektron” system [Complete]
JEM airlock relieve completion confirmation [Complete]
VIABLE-visual inspection of kits [Complete]
IMS update [Complete]
??? Maintenance
VIABLE – Payload photographing inside ISS [Complete]
SEISMOPROGNOZ. Data copying from ???? HDD to external media r/g 5701 [Complete]
DPC (S-band) [Complete]
On MCC Go ??? ?1 absorption cartridge regeneration (finish) [Complete]
SEISMOPROGNOZ. Archiving completion. r/g 5701 [Complete]
Task List

ARED Cylinder Evac
WHC KTO replace
Roscosmos Report
Ekon Photo
P/TV Lab CAMC swap [Completed Saturday]
ARED Dashpot inspect
CEVIS Isolator MOD
Cygnus pre-gather
EMU 3005 Photos
EMU LLB Auto-Init
EMU LLB SSC Relocate
EVA ORU Bag Inspection
IFM MOOT restow
IFM NOD1OS2 LHA R&R [Complete]
IFM NOD1PD2 LHA R&R [Complete]
NOD2O2 Cleanout
No-skid tape audit
NOD3 PEPS assessment
P/TV Hardware questions
PMM cleanup
WRS MRF Remove

Nominal commanding
JEM Airlock Monitoring/Leak Check
Three-Day Look Ahead:

Tuesday, 05/27: JEMAL Ops, CLPA install onto SSRMS B, Recycle tank R&R, Resist Tubule Sample removal
Wednesday, 05/28: 39S Launch and Dock, Sprint Ultrasound
Thursday, 05/29: Safety Briefing, CQ outfitting, PEPS inspection, SLM operations
QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – On
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Override
Carbon Dioxide Removal Assembly (CDRA) Node 3 – On
Major Constituent Analyzer (MCA) Lab – Stop
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Standby
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.